OpenStudy (anonymous):

tom makes $183.20 in 4 days. How much does he make in 3 days?

OpenStudy (anonymous):

Set up your equation: d(the number of days)*x(the variable in place for how much he makes a day)=m(money he makes) Plug in what you know 4x=183.20 Solve for x Then multiply x by three to find your answer.

OpenStudy (anonymous):

183.20/4 = x/3 cross multiply (4)(x) = (183.20)(3) 4x = 549.6 x = 549.60/4 x = 137.40 137.40 in 3 days

OpenStudy (anonymous):

@kelliegirl33 Please let them try it first before solving it for them. Code of conduct to the bottom left ;)

OpenStudy (anonymous):

should I delete it

OpenStudy (anonymous):

It's fine now. Just be helpful next time instead of giving it to them, they should work for it like we have right?

OpenStudy (anonymous):

yes...I will do better